Find the x-intercept and y-intercept for (x)
f(x)=x^2-8x+13

Answer by TimothyLamb(4379)

f(x) = x^2 - 8x + 13
---
the above quadratic equation is in standard form, with a=1, b=-8, and c=13
---
to solve the quadratic equation, plug this:
1 -8 13
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
x-intercepts:
the roots of the quadratic are real, therefore the x-intercepts are
( 5.73205081, 0 )
( 2.26794919, 0 )
---
y-intercept:
f(0) = 0^2 - 8(0) + 13 = 13
( 0, 13 )
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php